OpenStudy (anonymous):

Solve the absolute value inequality: /x+1/2/ - 1/2<3 1/2

OpenStudy (jdoe0001):

\(\bf \left|x+\cfrac{1}{2}\right|-\cfrac{1}{2}<3\cfrac{1}{2}\implies \left|x+\cfrac{1}{2}\right|<3\cfrac{1}{2}+\cfrac{1}{2}\\ \quad \\ \left|x+\cfrac{1}{2}\right|<3\cfrac{1}{2}+\cfrac{1}{2}\implies \begin{cases} +\left(x+\cfrac{1}{2}\right)<3\cfrac{1}{2}+\cfrac{1}{2}\\ \quad \\ \bf -\left(x+\cfrac{1}{2}\right)<3\cfrac{1}{2}+\cfrac{1}{2} \end{cases}\)

OpenStudy (jdoe0001):

what would you get for \(\bf 3\cfrac{1}{2}+\cfrac{1}{2}\quad ?\)

OpenStudy (anonymous):

uhh im not sure @jdoe0001

OpenStudy (jdoe0001):

well... if we were to make \(\bf 3\cfrac{1}{2}\) with only 1 numerator and 1 denominator.... what would it be?

OpenStudy (jdoe0001):

to make a "mixed fraction" a so-called an improper fraction you'd multiply the front number times the denominator plus the numerator all that over, the denominator left intact

OpenStudy (anonymous):

3 1/2 is 3.5 I know that.

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

rigth.... so.... ahemm \(\bf 3\cfrac{1}{2}\implies \cfrac{3\cdot 2+1}{2}\)

OpenStudy (jdoe0001):

so... what would that give you?

OpenStudy (anonymous):

3*2+1/2=6.5 ?

OpenStudy (jdoe0001):

well... \(\bf \cfrac{3\cdot 2+1}{2}\ne 3\cdot 2+\cfrac{1}{2}\)

OpenStudy (jdoe0001):

3.2 * 1/2 yes would be 6.5 but is not the same

OpenStudy (anonymous):

ahhh, ok

OpenStudy (jdoe0001):

woops, I meant 3*2 + 1/2 that is =)

OpenStudy (jdoe0001):

anyhow... ... what would you get for \(\bf 3\cfrac{1}{2}\implies \cfrac{3\cdot 2+1}{2}\quad ?\)

OpenStudy (anonymous):

:/ idk

OpenStudy (jdoe0001):

hmm notice ... \(\large \cfrac{3\cdot 2+1}{2}\) <--- is rather simple